# Inter-turbine Modelling of Wind-Farm Power using Multi-task Learning

### v1

@@ Abstract
This work first introduces a probabilistic regression model for predicting wind-turbine power, which adjusts for wake effects learnt from data. Spatial correlations in the learned model parameters for different tasks (turbines) are then leveraged in a hierarchical Bayesian model (an approach to multi-task learning) to develop a “metamodel”, which can be used to make power-predictions which adjust for turbine location—including on previously unobserved turbines not included in the training data. The results show that the metamodel is able to outperform a series of benchmark models, and demonstrates a novel strategy for making efficient use of data for inference in populations of structures, in particular where correlations exist in the variable(s) of interest (such as those from wind-turbine wake-effects).
@@ 1.2 Data-driven Modelling
A common application of data for online monitoring of wind-farms is in modelling the relationship between wind speed and power (known as power-curve modelling). These models may be used as a damage-sensitive feature during operation, to inform the need for any remedial repairs [4, 5]. As an additional benefit, these models can also be repurposed to make predictions of future wind-power generation (given forecasted weather inputs), which can help with risk mitigation in both operational and electricity market settings; this includes planning unit dispatch, maintenance scheduling and maximising profit in power trading [6, 7]. A relatively recent review of the modelling approaches developed for power-curve monitoring is provided by Wang et al. [8], which is a heavily researched area. However, in many cases the approaches used provide point, deterministic, predictions to predict output power. Inspection of typical power-curve data reveals that it is both heteroscedastic (the variance of the output power changes with wind speed), and asymmetrically distributed. Deterministic approaches fail to capture these properties; as such, probabilistic approaches have gained greater attention more recently from researchers [4, 5, 9], which provide information on both the mean and the expected variance of the prediction. This additional information gives greater insight into model uncertainty, which could be helpful for managing risk in both engineering and commercial settings.
@@ 1.3 Multi-task Learning
As a solution to the challenges of limited data, Transfer Learning (TL) is a machine-learning technique, which aims at improving the performance of models on target domains (tasks), by transferring knowledge from one or more related source tasks. One such flavour of TL, is multi-task learning (MTL), where the goal is to learn multiple related tasks simultaneously, so that similarities (and differences) can be harnessed between tasks; in this way, accuracy is also improved in domains where data are limited [17, 18, 19]. In practice, a common approach to MTL is by the use of hierarchical Bayesian models, which are able to pool information across tasks via population-level parameters, whilst accounting for task-specific nuances via task-level parameters. By pooling information in this way, these models can better handle tasks with limited data, and produce robust probabilistic predictions. This model architecture can intuitively be applied to wind-farms, where individual turbines (tasks) may have individual differences (e.g. because of maintenance, damage or physical location), but are ultimately similar to one another as part of the wind-farm population. Bayesian hierarchical models have recently been applied to other engineering infrastructure—Bull et al. [20] uses a hierarchical Bayesian modelling approach to learn population-level and task-specific parameters in a combined inference, for settings of a truck fleet survival analysis. Similarly, Dardeno et al. [21] and Brealy et al. [22] also use hierarchical Bayesian models, in order to transfer Frequency Response Function information between helicopter blades and wind-turbine foundation stiffness parameters respectively. In both cases, the uncertainty of parameters in data-sparse domains was reduced as a result of knowledge transfer between structures, as compared to modelling the data-sparse domains independently.
@@ 1.4 Research Contribution
In this work a probabilistic power-curve model is first developed, that considers the heteroscedastic and asymmetric nature of power-curve data, using data from an in-service wind-farm. This model utilises “global” wind speed and direction features, with predictions indicating that a level of wind-directionality (from wind-turbine wake effects), has been captured by the model. Spatial correlations found in the underlying model parameters inspired the incorporation of this model into a partially-pooled hierarchical Bayesian model architecture, where the population-level parameters were designed to capture spatial inter-task correlations between turbines, which is described here as the “metamodel”.
{mdframed}
This model design has a tangible real-world benefit - if one were able to make accurate probabilistic predictions for a variable of interest, at a location without data (and perhaps even without associated telemetry), this may reduce the overall telemetry requirements among a population.
The results show that for predicting wind-turbine output power, the metamodel outperforms a series of benchmark models, both in terms of mean prediction and predictive uncertainty, making efficient use of data and accounting for environmental variability. The metamodel could also be described as a purely data-based approach to power modelling with adjustment for wake effects, without the need for physics-based simulation (thus making it more efficient).
@@ 2.1 Description
A Supervisory Control and Data Acquisition (SCADA) dataset that has recently become available to the authors of this work, covers 224 wind-turbines, spanning three separate wind-farms between the years 2020-2023, at a ten-minute resolution. The measurements within come from a broad range of sensors, but generally fall into electrical, control, component/system temperature measurements, and ambient environmental parameters. This study focuses on one of these wind-farms (under the pseudonym “Ciabatta”), for the full year of 2021; this was chosen to maintain a full year of cyclical environmental variation, and to reduce computational overhead with a focus on demonstrating the methods herein.
The raw SCADA data, showing the power-curve relationship for the wind-turbines in wind-farm Ciabatta are shown in blue in Figure 1. In these data, a number of patterns typical to wind-farm operation are identifiable, including curtailments (whereby turbine power is deliberately limited), shutdowns (indicated by a power output of zero, despite significant wind speeds), and power-boosting (where power is raised above nominal turbine capacity). These three (controlled or otherwise) deviations of power away from normal operation are typically related to technical and/or commercial constraints, dependent on factors which were considered outside the scope of this work.
@@ 2.2 Data Filtering
For the purposes of predicting power output during normal operation, data most likely related to shutdowns and curtailments were filtered out. For the power-boosted data, since for some turbines this made up a large majority of the data at the associated wind speeds, during filtering, model training and testing this was artificially reduced to rated power; this was done to avoid creating data-sparse regions for some turbines, which would have made model comparison more difficult. This pre-processing was also considered reasonable from a process understanding, since wind-turbine wake effects (discussed in the following section) are more prevalent at lower wind speeds [23].
While filtering the data was a significant undertaking in itself, a summary of the main criteria chosen is shown in Table 1.
General approach Features Filtering purpose
@@
thresholding blade pitch angle and RPM curtailments and shutdowns
rate of change RPM transient startup and shutdown procedures
time stationarity power and freestream wind speed curtailments and erroneous data
@@
Mahalanobis squared-distance thresholding [24] blade pitch angle vs. power outliers
@@
standard deviation thresholding freestream wind speed vs. power outliers
Table 1: Summary of the approaches used to filter the raw SCADA data
The filtered data are plotted in orange in Figure 1, compared to the raw unfiltered data in blue. It can be seen that the filtered data are much more representative of normal operation only. Considering these filtered data (and power-curve data generally), there are two distinct features that present challenges for modelling: (1) the data are heteroscedastic—as the wind speed changes, the variance changes significantly and (2), the skewness changes from being positive at low wind speeds before cut-in, to negative above the rated wind speed. These properties should be considered in model design to ensure the data are modelled appropriately.
@@
Figure 1: Data from all turbines in wind-farm Ciabatta showing the power-curve relationship. Blue markers show the data prior to filtering, and orange markers show the data after filtering.
@@ 2.3 Model Training Features
For all models developed in this study, three features were used for predicting turbine-specific output power. First, came the freestream wind speed (later referred to as Feature One); this was calculated as the maximum measured wind speed for all turbines used for training, for a given timestamp, since data from a meteorological mast was not available. The analysis also assumed equivalence between the turbine measured yaw angles and actual wind direction, and took the sine and the cosine of the yaw angle; this was done to create a smooth transition between the equivalent angles of 0 and 360 degrees. Median values were taken for these sine and cosine features for a given timestamp (and are later referred to as Features Two and Three), since there were discrepancies between individual turbine yaw angles. Finally, all features were scaled between zero and one for both numerical efficiency, and data anonymity.
The motivations for choosing these features were as follows: (1) using the freestream wind speed in combination with wind directional features, enables the possibility to capture some of the wake effects that may be present in the data. (2) While not the main focus of this study, population-level features are more representative of forecast data by data providers such as ECMWF [25], since higher resolution turbine-specific forecasts that also account for turbine wake effects are not readily available.
@@ 3 B-spline Beta Regression Model
The heteroscedacity, skewness and bounded nature of power-curve data share similarities with the shape of the beta distribution, which is bounded between zero and one, and whose variance changes with the mean of the distribution. These two factors motivated the development of the B-Spline Beta Regression Model (BSBR), which is a probabilistic regression-based model utilising a beta-likelihood. Mclean et al. [4] similarly use a beta-likelihood, and the description of the BSBR model that follows is very similar to that of Capelleti et al. [26] whom also used a spline-based approach.
@@ 3.2 Beta Regression Model
Beta regression is one such type of GLM, where the beta distribution is chosen for modelling $\mathbf{y}$. Beta distributions are bounded in the interval between zero and one, whilst their variance can change with the mean of the distribution; this has clear similarities with power-curve data, which can be sensibly bounded between zero and rated power, with trivial rescaling. The beta distribution for a scalar $y$ is defined as,
@@
where $\Gamma(\cdot)$ denotes the gamma function, and where $10<\mu<$ 1 and $\phi>0$. The mean and variance of $y$ are,
@@
In this case for modelling power-curve data, first consider a linear predictor model, $f_{1}=\mathbf{X}\boldsymbol{\beta}$ for the GLM, where $\boldsymbol{\beta}$ is a vector of regression parameters. For computational efficiency during Markov Chain Monte Carlo (MCMC) sampling (the method for parameter estimation–see Section 5.2) it is better to use the QR reparameterisation such that $\mathbf{X}=\mathbf{QR}$ [27]. If one also lets $\boldsymbol{\eta}=\mathbf{R}\boldsymbol{\beta}$ then one has $f_{1}=\mathbf{Q}\boldsymbol{\eta}$. For the final component of the GLM - the link function is chosen to be the logit function (equivalent to $g^{-1}$ equalling the expit function [28]), such that,
$\boldsymbol{\mu}=expit(\mathbf{Q}\boldsymbol{\eta})$ (4)
where $\boldsymbol{\eta}$ is a vector of regression parameters (a transformed version of $\boldsymbol{\beta}$), and $\boldsymbol{\mu}$ is the mean parameter vector. The choice of the logit link function ensures that $10<\mathbb{E}(\mathbf{y})<$ 1, which in words, means the predictive mean of the model must lie between zero and one, intuitively matching the behaviour of (normalised) power-curve data. For additional flexibility in the model, the precision parameter vector $\boldsymbol{\phi}$ was also allowed to vary with $\mathbf{X}$ according to an additional linear model $f_{2}=\mathbf{X}\boldsymbol{\theta}$, where $\boldsymbol{\theta}$ is an additional vector of regression parameters. Similarly using the QR reparameterisation, one can write $f_{2}=\mathbf{Q}\boldsymbol{\zeta}$, where $\boldsymbol{\zeta}=\mathbf{R}\boldsymbol{\theta}$. Since $\boldsymbol{\phi}$ must be non-negative (to ensure that $\mathbb{V}(\mathbf{y}$) is positive), a natural logarithm link function is chosen (equivalent to $g^{-1}$ equalling the exponential function), such that,
$\boldsymbol{\phi}=exp(\mathbf{Q}\boldsymbol{\zeta})$ (5)
where $\boldsymbol{\zeta}$ is another vector of regression parameters (transformed from $\boldsymbol{\theta}$). This model definition allows both the mean, $\boldsymbol{\mu}$, and dispersion, $\boldsymbol{\phi}$, parameter vector entries to vary as a function of $\mathbf{X}$ via $\mathbf{Q}$, resulting in beta distributions whose shapes vary with the input feature(s); this is similar to the “Variable dispersion beta regression model” described in [26].

### v2

@@ Abstract
This work first introduces a probabilistic regression model for predicting wind-turbine power, which adjusts for wake-effects learned from data. Spatial correlations in the learned model parameters for different tasks (turbines) are then leveraged in a hierarchical Bayesian model (an approach to multi-task learning) to develop a “metamodel”, which can be used to make power-predictions which adjust for turbine location—including on previously unobserved turbines not included in the training data. The results show that the metamodel is able to outperform a series of benchmark models, while demonstrating a novel strategy for making efficient use of data for inference in populations of structures, in particular where correlations exist in the variable(s) of interest (such as those from wind-turbine wake-effects).
@@ 1.2 Data-driven Modelling
A common application of data for the online monitoring of wind-farms is in modelling the relationship between wind-speed and power (known as power-curve modelling). These models may be used as a damage-sensitive feature during operation, to inform the need for any remedial repairs [4, 5]. As an additional benefit, these models can also be repurposed to make predictions of future wind-power generation (given forecasted weather inputs), which can help with risk mitigation in both operational and electricity market settings; this includes planning unit dispatch, maintenance scheduling and maximising profit in power trading [6, 7]. A relatively recent review of the modelling approaches developed for power-curve monitoring is provided by Wang et al. [8], which is a heavily researched area. However, in many cases the approaches used provide point, deterministic, predictions to predict output power. Inspection of typical power-curve data reveals that it is both heteroscedastic (the variance of the output power changes with wind-speed), and asymmetrically distributed. Deterministic approaches fail to capture these properties; as such, probabilistic approaches have gained greater attention more recently from researchers [4, 5, 9], which provide information on both the mean and the expected variance of the prediction. This additional information gives greater insight into model uncertainty, which could be helpful for managing risk in both engineering and commercial settings.
@@ 1.3 Multi-task Learning
As a solution to the problem of limited data, multi-task learning (MTL)—a form of transfer learning (TL)—is a machine-learning approach where the goal is to learn multiple related tasks simultaneously, so that similarities (and differences) can be harnessed between tasks; in this way, model predictive accuracy is improved in domains where data are limited [17, 18, 19]. In practice, a common approach to MTL is by the use of hierarchical Bayesian models, which are able to pool information across tasks via population-level parameters, while accounting for task-specific nuances via task-level parameters. By pooling information in this way, these models can better handle tasks with limited data, and produce robust probabilistic predictions. This model architecture can intuitively be applied to wind-farms, where individual turbines (tasks) may have individual differences (e.g. because of maintenance, damage or physical location), but are ultimately similar to each other as part of the wind-farm population. Hierarchical Bayesian models have recently been applied to other engineering infrastructure; Bull et al. [20] uses a hierarchical Bayesian modelling approach to learn population-level and task-specific parameters in a combined inference, in the setting of a truck-fleet survival analysis. Similarly, Dardeno et al. [21] and Brealy et al. [22] use hierarchical Bayesian models, in order to transfer Frequency Response Function information between helicopter blades and wind-turbine foundation stiffness parameters respectively. In both cases, the uncertainty of parameters in data-sparse domains was reduced as a result of knowledge transfer between structures, as compared to modelling the data-sparse domains independently.
@@ 1.4 Research Contribution
In this work, a probabilistic power-curve model is first developed, that considers the heteroscedastic and asymmetric nature of power-curve data, using data from an in-service wind-farm. This model utilises wind-farm (population-level) wind-speed and direction features, with predictions indicating that a level of wind-directionality (from wind-turbine wake-effects), has been captured by the model. Spatial correlations found in the underlying model parameters inspired the incorporation of this model into a hierarchical Bayesian model architecture, where the population-level parameters were designed to capture spatial, inter-task correlations between turbines, which is described here as the “metamodel”.
This model design has a tangible real-world benefit—if one were able to make accurate probabilistic predictions for a variable of interest, at a location without data (and perhaps even without associated telemetry), this may reduce the overall telemetry requirements among a population.
The results show that for predicting wind-turbine output power, the metamodel outperforms a series of benchmark models, both in terms of mean prediction and predictive uncertainty, by making efficient use of data and accounting for environmental variability. The metamodel could also be described as a purely data-based approach to power modelling with adjustment for wake-effects, without the need for physics-based simulation (thus making it more efficient).
@@ 2.1 Description
A Supervisory Control and Data Acquisition (SCADA) dataset was available to the authors of this work, which contains 224 wind-turbines, spanning three separate wind-farms between the years 2020-2023, at a ten-minute resolution. The measurements within come from a broad range of sensors, but generally fall into electrical, control, component/system temperature measurements, and ambient environmental parameters. This study focuses on one of these wind-farms (under the pseudonym “Ciabatta”), for the full year of 2021; this was chosen to maintain a full year of cyclical environmental variation, and to reduce computational overhead with a focus on demonstrating the methods herein.
The raw SCADA data, showing the power-curve relationship for the wind-turbines in wind-farm Ciabatta are shown in blue in Figure 1. In these data, a number of patterns typical to wind-farm operation are identifiable, including curtailments (whereby turbine power is deliberately limited), shutdowns (indicated by a power output of zero, despite significant wind-speeds), and power-boosting (where power is raised above nominal turbine capacity). These three (controlled or otherwise) deviations of power away from normal operation are typically related to technical and/or commercial constraints, dependent on factors which were considered outside the scope of this work.
@@ 2.2 Data Filtering
For the purposes of predicting power output during normal operation, data most likely related to shutdowns and curtailments were filtered out. For the power-boosted data, rather than filtering it, it was instead artificially reduced to the rated-power; this was done for two main reasons: (1) For some turbines, it accounted for a significant portion of the data at high wind-speeds, such that filtering it would create data-sparse regions for those turbines, making robust model development and comparison between different models more difficult. (2) From a practical perspective, if a model predicts that rated-power is achievable, and the specific conditions required for power-boosting are known, then domain-experts could apply a manual correction if necessary. Additionally, since wake-effects are more prevalent at lower wind-speeds [23], this modification was not expected to significantly interfere with wake-related behaviour in the data. A summary of the main criteria chosen to filter the data are shown in Table 1.
Filtering step Description Filtering target
@@
1: Thresholding Filter data based on blade pitch angle and rotor speed (RPM) Curtailments and shutdowns
2: Rate of change Filter data where the RPM changes rapidly between time steps Transient behaviour during startup and shutdown procedures
3: Stationary data Filter data where the power or wind-speed measurements are stationary Curtailments and sensor faults
@@
4: Mahalanobis squared-distance outlier detection [24] Filter data based on the distance in the blade pitch-angle vs. power relationship Remaining outliers
@@
5: Standard-deviation outlier detection Filter that retains the middle two standard deviations in power, for the binned freestream wind-speed Remaining outliers
Table 1: Summary of the steps used to filter the raw SCADA data
Between low and rated-power, in normal operation the blade pitch angle is approximately zero, while during curtailment it increases to control the output power to a set level. When a turbine is offline or during a shutdown, the pitch angle is also higher than in normal operation. Additionally, as the wind-speed increases beyond the cut-out wind-speed, the shaft speed (RPM) remains high, while the generated power drops considerably; these three sets of behaviours were targeted for filtering in step one, by defining set-thresholds on the blade-pitch-angle and rotor-speed, for differing levels of power. The second filtering step was introduced as it was found that in some cases, the RPM changed quickly; this was considered indicative of startup and shutdown procedures of the turbines, and not representative of normal operation. In this step, data where the RPM increased or decreased by more than a set threshold, compared to the previous value were filtered out. In the data, instances of stationary wind-speed measurements were found, which were deemed likely to be due to sensor or SCADA system error. To combat this, step three was used to remove data where the wind-speed measurements were stationary in time. The same technique was also applied to remove any stationary power data that was not filtered by step one. Step four was used to help remove any remaining outliers in the pitch-angle vs. power relationship, by using a Mahalanobis Squared Distance (MSD)-based approach [24], which takes account of both correlation and scale in the data. Finally, step five removed remaining outliers that were observed in the output power, by retaining only the middle two standard-deviations for a binned freestream wind-speed—defined in Section 2.3.
The filtered data are plotted in orange in Figure 1, compared to the raw unfiltered data in blue. It can be seen that the filtered data appear much more representative of normal operation only.
@@
Figure 1: Data from all turbines in wind-farm Ciabatta showing the power-curve relationship. Blue markers show the raw data prior to filtering, and orange markers show the data after filtering.
To demonstrate how the filtering process influences the underlying distributions in the data, Figure 2 shows density plots for the generated power (2(a)), yaw angle (2(b)), and wind-speed (2(c)) variables. The raw (unfiltered) data are shown in blue and the filtered data are shown in orange. Here it can be seen that in most cases the general shapes of the distributions after filtering are consistent. One clear difference can be seen at low (or zero) generated power; this is expected, since a large amount of data was present in this region and targeted for filtering. Some reductions in data density can be seen between 0.4–0.6 in generated power; this is because of the removal of the curtailment data. Finally, there is an increase in the quantity of data in the final bin of the filtered generated power; this is because of the capping of the power-boosted data. For completeness, the filtered data that had then undergone stratified (sub)sampling on both yaw angle and power features (discussed in Section 5.1.2) are also shown here in green; in this case it can be seen that the desired effect of flattening the density plots for the yaw angle and power features had broadly been achieved.
(a) Generated Power.
(b) Yaw angle.
(c) Wind-speed.
Figure 2: Density plots for the raw data (blue), filtered data (orange) and the subsampled data (green) for the generated power, yaw angle and wind-speed variables.
Considering these filtered data (and power-curve data generally), there are two distinct features that present challenges for modelling: (1) the data are heteroscedastic—as the wind-speed changes, the variance in generated power changes significantly and (2), the skewness changes from being positive at low wind-speeds before cut-in, to negative above the rated wind-speed. These properties should be considered in the model design phase to ensure that the data are modelled appropriately.
@@ 2.3 Model Training Features
For all models developed in this study, three features were used for predicting turbine-specific output power. First, came the freestream wind-speed (later referred to as Feature One); since data from a meteorological mast was not available, this was instead approximated as being equivalent to the maximum measured wind-speed of the turbines used in training, for a given timestamp. Since there are turbines on all edges of the wind-farm in the training data sets, this was deemed a reasonable approximation as measurements from “unwaked” turbines were always available. The authors are aware that anemometer sensor error will likely add noise to the data and could reduce model predictive performance as a result, however it is believed it does not detract from the novelty of the method developed in this work.
The second and third features were derived from what was considered a proxy for wind-direction—the measured turbine yaw angle; these were assumed to be approximately equal to each other, since for operation in the normal regime the control system should orientate the turbines in the direction of the incoming wind. Specifically, Features Two and Three took the sine and the cosine of the yaw angle; this was done to create a smooth transition in the inputs between the equivalent angles of zero and 360 degrees. The median values were used for these features which were taken across the turbines for a given timestamp, to obtain an averaged wind-direction, less subject to individual turbine noise. Finally, all features were scaled between zero and one for both numerical efficiency, and data anonymity.
The motivations for choosing these features were as follows: (1) using the freestream wind-speed in combination with wind-directional features, enables the possibility to capture some of the wake-effects that may be present in the data. (2) While not the main focus of this study, population-level features are more representative of forecast data by data providers such as ECMWF [25], since higher resolution turbine-specific forecasts that also account for turbine wake-effects are not readily available.
@@ 3 B-spline Beta Regression Model
The heteroscedacity, skewness and bounded nature of power-curve data share similarities with the shape of the beta distribution, which is bounded between zero and one, and whose variance changes with the mean of the distribution. These two factors motivated the development of the B-Spline Beta Regression Model (BSBR), which is a probabilistic regression-based model utilising a beta-likelihood. Mclean et al. [4] similarly used a beta-likelihood, and the description of the BSBR model that follows is very similar to that of Capelletti et al. [26] whom also used a spline-based approach.
@@ 3.2 Beta Regression Model
Beta regression is one such type of GLM, where the beta distribution is chosen for modelling $\mathbf{y}$. Beta distributions are bounded in the interval between zero and one, while their variance can change with the mean of the distribution; this has clear similarities with power-curve data, which can be sensibly bounded between zero and rated power, with trivial rescaling. The beta distribution for a scalar $y$ is defined as,
@@
where $\Gamma(\cdot)$ denotes the gamma function, and where $0<\mu<1$ and $\phi>0$. The mean and variance of $y$ are,
@@
In this case for modelling power-curve data, consider a linear predictor model, $f_{1}=\mathbf{X}\boldsymbol{\eta}$ for a GLM, where $\boldsymbol{\eta}$ is a vector of regression parameters. The link function is chosen to be the logit function (equivalent to $g^{-1}$ equalling the expit function [27]), such that,
$\boldsymbol{\mu}=expit(\mathbf{X}\boldsymbol{\eta})$ (4)
where $\boldsymbol{\mu}$ is the mean parameter vector for the beta distribution. The choice of the logit link function ensures that $0<\mathbb{E}(\mathbf{y})<1$, which in words, means the predictive mean of the model must lie between zero and one, intuitively matching the behaviour of normalised power-curve data. For additional flexibility in the model, the precision parameter vector $\boldsymbol{\phi}$ (in the beta distribution) was also allowed to vary with $\mathbf{X}$ according to an additional linear model $f_{2}=\mathbf{X}\boldsymbol{\zeta}$, where $\boldsymbol{\zeta}$ is an additional vector of regression parameters. Since $\boldsymbol{\zeta}$ must be non-negative (to ensure that $\mathbb{V}(\mathbf{y}$) is positive), a natural logarithm link function is chosen (equivalent to $g^{-1}$ equalling the exponential function), such that,
$\boldsymbol{\phi}=exp(\mathbf{X}\boldsymbol{\zeta})$ (5)
This model definition allows both the mean, $\boldsymbol{\mu}$, and precision, $\boldsymbol{\phi}$, parameter vector entries to vary as a function of $\mathbf{X}$, resulting in beta distributions whose shapes vary with the input feature(s); this is similar to the “Variable dispersion beta regression model” described in [26].